# Richelieu: Self-Evolving LLM-Based Agents for AI Diplomacy

### v1

@@ 1 Introduction
Diplomacy, as a cornerstone of international relations, is an intricate and multifaceted activity that lies at the heart of human society’s most complex interactions. It encompasses a wide array of skills and strategies, including social reasoning, negotiation, and long-term planning, to navigate the intricate web of relationships and alliances between multiple parties. Mirroring this complexity, the board game Diplomacy[59] involves seven players to control European powers, presenting a complex strategic challenge that requires both sophisticated negotiation and strategic planning to triumph.
@@
The AI community has shown an increasing interest in the deployment of AI agents to master such games [45, 26, 29, 15, 36, 28]. The recent breakthrough [6] has turned into press diplomacy, which allows communication between players. However, the previous methods [6] heavily rely on domain-specific human data, leading to its poor generalization to other scenarios/ applications. The question then arises: Can we build an AI agent that excels in the art of diplomacy without relying on domain-specific human data?
@@
Recently, agents based on the Large Language Model(LLM) have emerged as a promising development for AI agents. The previous applications on personal assistants [32], robotics [10, 64], and video games [48] have shown the surprising ability of LLM-based agents in communication and planning, benefiting from the emergent ability of common sense reasoning, in-context/ few-shot learning, and sophisticated natural language processing on LLMs. However, diplomacy presents a unique set of challenges. It not only requires planning long-horizon strategic [40] and communicating with natural language, but also reasoning and adopting the complex social dynamics with partial observations, including gaining trust and reputation, building rapport, detecting deception, and assessing the reliability of other players.
@@
In this work, we aim to make the first attempt to explore LLMs’ potential to develop a human-like AI diplomacy agent. We name the agent Richelieu in memorizing a pivotal figure in European history who had enduring impacts on French politics, foreign affairs, and state building. To achieve this goal, we have identified three core and essential capabilities that are crucial for building an LLM-based societal agent: 1) Social reasoning This is the basic function for a social agent to interact with others, particularly for adapting to the dynamic changes in the nation’s intentions and relationships. 2) Balance long- and short-term planning Diplomacy often requires a delicate balance between short-term tactics and long-term strategies. An effective AI agent must be able to recognize and weigh the immediate consequences of its actions against the potential long-term outcomes. 3) Powerful memory A robust memory system is a critical component of learning and improvement. The AI agent must be able to recall and integrate information from past negotiations and actions to inform its current and future decision-making processes. This endows the agent with the ability to evolve. 4) Profound reflection An AI agent capable of profound reflection can analyze its own decisions, learn from its memory experience, and adapt its strategies accordingly. By integrating these three capabilities, the agent can operate at the highest level of diplomatic sophistication, outperforming the state-of-the-art AI diplomats [6].
@@
Our contributions can be summarized in three-fold: 1) We introduced a new paradigm for building AI diplomacy agents, compared to previous work (Fig. 1). The agent can self-evolve by generating experience via self-play games, without any task-specific human data. 2) We demonstrate the superior performance of our agent playing against the SOTA method, e.g., Cicero [6], that relies on a large-scale human demonstration for training. 3) We further analyze the effectiveness of each module in our agent and the generalization of our agent in adopting different LLMs, such as GPT4.0 and Llamma 3.
@@ 2 Related work
AI Diplomacy. The game involves seven players controlling different powers in Europe. In each turn, players can negotiate for cooperation before making moves to take as much supply centers as they can. Apparently, this challenging strategy task requires both complex negotiation skills and superior planning capability for player agents to achieve final victory. So far, most AI research on this task remain focused on the planning strategies (a.k.a. No-Press Diplomacy where no communication channels are allowed). The setting remains challenging considering its enormous action space of $110^{2}$ 1 to $410^{6}$ 4 per turn (compared with Chess, which has much fewer than 100 actions per turn). No wonder existing efforts rely on human data to play the game. Among the methods, one typical research is DipNet [39] which uses supervised and reinforcement learning. Based on DipNet, BRPI [3], SearchBot [18], DORA [5], and KL-Regularized search (Diplodocus) [24] were conducted. Until very recently, research has also emerged for the full-setting of Diplomacy, or Press Diplomacy where players are allowed to communicate with each before making their moves in each turn. Such studies [13][6][25][29] mainly benefit from the recent thriving language models. Specifically, notable advancements include policy iteration methods from DeepMind and Facebook AI Research’s equilibrium search agent [25]. However, Deepmind propose to learn negotiation agents based on predefined contracts/protocols [29]. And Meta AI’s work, instead of one unified architecture, Cicero [6] integrates a language model for negotiation and an RL model for planning respectively. Such separately trained models make it inconvenient for agents’ continual evolution. What’s more, like no-press methods, these approaches heavily rely on human player data for agent training. Unlike these approaches, this paper delves into solving the negotiation and planning in one single self-evolving LLM-based agent model, without any pre-collected human expert training data.
@@
LLM-based Agents. With the emergence and growth of large language models (LLM), there is a growing trend in utilizing LLMs as fundamental controllers for autonomous agents[52]. One wide application genre is LLM-based answering engines, which merely cover the negotiation aspects of Diplomacy. Such systems include HuggingGPT [44], GPT4Tools [63] and ToT [65], etc. They leverage LLMs to manage Al models, use tools, implement policy iteration, and enhance problem-solving across various tasks. Related work including AutoGPT, AgentGPT, BabyAGl [47], Toolformer [43], and Visual ChatGPT aim to improve LLM capabilities in task automation and tool usage. And Reflexion, a framework that improves LLMs through linguistic feedback and episodic memory [68], facilitating better decision-making across diverse tasks is proposed. Besides [55][50][56][76][62] apply LLM agents to the complex planning tasks in the well-known open-world game Minecraft[16]. Unlike these LLM-based agents which only focus on the negotiation/planning aspect, the proposed approach involves multiple self-evolving schemes to handle both of them simultaneously.
@@ 3 Problem Statement
The Diplomacy game [59, 8] is set in pre-World War I Europe and involves each player (agent) representing one of the seven Great Powers of Europe, such as Germany, France, England, Italy, Austria-Hungary, Russia, and Turkey. Each player has a set of military units, including armies and fleets, which they can move and use to capture other supply centers. The ultimate goal for the agent is to control a majority of the total supply centers on the board by the end of the game’s Fall phase. It’s important to note that it is not won by eliminating other players or their units; it is won by controlling the requisite number of supply centers. This often involves forming and breaking alliances, negotiating, and sometimes betraying other players to achieve one’s own goals.
In each turn,[58, 53] the agent $i$ gets the current state $s_{t}\in S$, the actions of other players from the previous turn $\vec{a}^{-i}_{t-1}$, and the messages $\vec{m}^{-i,i}_{t}$ from other players during this turn’s negotiations. The state $s_{t}$ for the environment includes the ownership of each territory on the map by a particular country and where the armies of each country are located. Based on this information, the agent needs to engage in negotiations with other players, sending messages $\vec{m}^{i,-i}_{t}$ to chat with other players, and then take the actions $a^{i}_{t}$ in this turn. The possible actions an agent can take $a^{i}_{t}\in A$ are commands to the armies, such as moving into an adjacent territory, supporting another unit, or holding a position. Actions can also include diplomatic moves, such as proposing or withdrawing from an alliance, although these are less formalized in the game mechanics.[39, 22]
@@ 4 Self-Evolving LLM-based Diplomat
We have constructed a comprehensive framework with modules for memory management, social reasoning, strategic planning, negotiation, decision-making, memory update, and self-evolving to fully leverage the capabilities of LLMs. Richelieu starts by setting up with map details, game rules, domain knowledge, and the long-term goal.[74, 58, 53] At each turn, the agent will run in the following steps: 1) Social Reasoning: First of all, the agent undergoes a comprehensive analysis of the game state $s_{t}$ to build the social belief, including the intention of other players and their relationship $\vec{\phi_{t}}\in\Phi^{n}$.[71, 19] 2) Planner with Reflection: Then, the agent proposes sub-goals $\chi^{i}_{t}\in X$ that is strategically aligned with the long-term goals $\Upsilon$, with the social belief and refining the proposed goal with experience $\vec{\eta_{t}}\in H^{m}$ abstract from the memory $M$ via self-reflection.[51, 57] 3) Negotiator: To achieve the sub-goals, the negotiator will start a dialogue session with some players, and evaluate their trueness $\vec{\psi}^{-i}_{t}$ by referring to their words $\vec{m}^{-i,i}_{t}$, the current state $s_{t}$, their sincerity $\vec{\gamma}^{-i}_{t}$ and the experience $\vec{\xi_{t}}$[1, 7] 4) Actor: After negotiation, the actor decides its course of action $a^{i}_{t}$, based on the sub-goal $\chi^{i}_{t}$ and updated social state $s_{t+1}$, marking the end of that turn. 5) Memory Management: The state of the current turn $s_{t}$, the content of negotiations $\vec{m_{t}}$, the actions taken by all players $\vec{a_{t}}\in A^{n}$, and the sub-goals set forth $\chi^{i}_{t}$ are all logged within the memory as $\mu\in M$. This logged data serves as a historical experience, guiding Richelieu’s subsequent actions in future turns [20, 73]. 6) Self-evolution: The agent’s evolution is highly dependent on the diversity of experiences stored in its memory. As this diversity grows, so does the agent’s capability. Without human demonstrations, we employ multi-agent self-play games, i.e., our agents respectively control all the countries to simulate and acquire diverse experiences for self-evolving. Notably, the agent can further evolve during testing to adapt to different players.
@@ 4.1 Social Reasoning
There are no permanent enemies, no permanent allies. The relationship among countries is dynamically changing upon the evolving global state. However, it is difficult to determine the appropriate allies and enemies with partial observation. For example, there is uncertainty about the intentions of potential allies, which could lead to betrayal at pivotal moments. Consequently, we need to identify the intention and relationship of the current state by social reasoning to shape the social belief [71, 19].
1) Modeling Relationship: Before setting sub-goals, Richelieu evaluates its relations with other players, identifying enemies such as aggressive nations, vulnerable neighbors for expansion, and those with long-term potential threats. It also seeks out potential allies to counter these threats.[46, 72] Simultaneously, Richelieu also tries to identify potential allies that could be instrumental in countering these adversaries. By isolating the analysis of inter-player relationships as a discrete element, Richelieu strategically exploits the actions of other players in subsequent stages of the game to reach its goals. 2) Inferring Intention: The social belief is also used by the planner, ensuring that Richelieu’s sub-goals are formulated with a comprehensive consideration of the behaviors and intentions of other intelligent agents within the game.[14, 21] Richelieu’s sub-goals will particularly emphasize on those who are identified as potential adversaries or allies, fostering more effective collaboration with potential allies and participation in strategic opposition against adversaries. Furthermore, the insights gleaned from this analysis are instrumental in the subsequent negotiation phases. They are employed to assess the authenticity of the statements made by other players, as well as to aid Richelieu in reaching cooperative agreements.
@@ 4.2 Strategic Planner with Reflection
The strategic planner specifies the sub-goals, which serves as an intermediary between immediate actions and the overarching goal of securing victory in the game. That is because we observe that LLMs are often characterized by their propensity to prioritize short-term gains in decision-making processes, with a notable deficiency in incorporating the future into their strategic calculations. [41, 70]For example, it is common for a non-neighboring country to become too powerful. Formally, $\vec{\chi_{t}}\leftarrow SR(s_{t},\vec{\phi_{t}},\Upsilon)$ where $\vec{\chi_{t}}=(\chi^{i}_{t},\chi^{1}_{t},\ldots,\chi^{n}_{t})$ represents the proposed sub-goals and other players’ intention that we inferred, $\vec{\phi_{t}}\in\Phi^{n}$ represents the inferred relationship on the social belief. These goals may encompass a range of tactical considerations, such as the containment of a formidable rival’s advancement or the strategic expansion in a particular direction to consolidate power.
Reflection with Memory. We further develop a reflection mechanism to enhance the rationality and effectiveness of our agent’s sub-goals in achieving long-term goals.[33] This reflection mechanism relies on the past experiences to critique and enhance proposed sub-goals. We employ a similarity-based function to find relevant historical experiences that match the current game state from its memory. This function considers two factors: goal similarity and state similarity, to select the most comparable experiences. The process can be written as: $\vec{\eta_{t}}\leftarrow h(s_{t},\chi^{i}_{t},M)$, where $\vec{\eta_{t}}\in H^{m}$. In practice, considering the limited context windows of LLM, we retrieve the most analogous experiences from the memory based on these metrics. Experiences with high evaluative scores reinforce successful strategies and support the continuity of existing sub-goals. On the other hand, lower scores indicate areas that need improvement and prompt the necessary adjustments. As our agent, Richelieu, undergoes more training sessions, its reflection abilities improve. The growing pool of historical experiences consistently enhances its performance.

### v2

@@ 1 Introduction
Diplomacy, as a cornerstone of international relations, is an intricate and multifaceted activity that lies at the heart of human society’s most complex interactions. It encompasses a wide array of skills and strategies, including social reasoning, negotiation, and long-term planning, to navigate the intricate web of relationships and alliances between multiple parties. Mirroring this complexity, the board game DiplomacyWikipedia (2024) involves seven players to control European powers, presenting a complex strategic challenge that requires both sophisticated negotiation and strategic planning to triumph.
@@
The AI community has shown an increasing interest in the deployment of AI agents to master such games (Shoker et al., 2023; Konya et al., 2023; Kramár et al., 2022; Duéñez-Guzmán et al., 2023; Mukobi et al., 2023; Kovač et al., 2023). The recent breakthrough (Bakhtin et al., 2022) has turned into press diplomacy, which allows communication between players. However, the previous methods (Bakhtin et al., 2022) heavily rely on domain-specific human data, leading to its poor generalization to other scenarios/ applications. The question then arises: Can we build an AI agent that excels in the art of diplomacy without relying on domain-specific human data?
@@
Recently, agents based on the Large Language Model(LLM) have emerged as a promising development for AI agents. The previous applications on personal assistants Li et al. (2024b), robotics Cheng et al. (2024); Yang et al. (2023c), and video games Wan et al. (2024) have shown the surprising ability of LLM-based agents in communication and planning, benefiting from the emergent ability of common sense reasoning, in-context/ few-shot learning, and sophisticated natural language processing on LLMs. However, diplomacy presents a unique set of challenges. It not only requires planning long-horizon strategic (Qi et al., 2024) and communicating with natural language, but also reasoning and adopting the complex social dynamics with partial observations, including gaining trust and reputation, building rapport, detecting deception, and assessing the reliability of other players.
@@
In this work, we aim to make the first attempt to explore LLMs’ potential to develop a human-like AI diplomacy agent. We name the agent Richelieu in memorizing a pivotal figure in European history who had enduring impacts on French politics, foreign affairs, and state building. To achieve this goal, we have identified three core and essential capabilities that are crucial for building an LLM-based societal agent: 1) Social reasoning This is the basic function for a social agent to interact with others, particularly for adapting to the dynamic changes in the nation’s intentions and relationships. 2) Balance long- and short-term planning Diplomacy often requires a delicate balance between short-term tactics and long-term strategies. An effective AI agent must be able to recognize and weigh the immediate consequences of its actions against the potential long-term outcomes. 3) Powerful memory A robust memory system is a critical component of learning and improvement. The AI agent must be able to recall and integrate information from past negotiations and actions to inform its current and future decision-making processes. This endows the agent with the ability to evolve. 4) Profound reflection An AI agent capable of profound reflection can analyze its own decisions, learn from its memory experience, and adapt its strategies accordingly. By integrating these three capabilities, the agent can operate at the highest level of diplomatic sophistication, outperforming the state-of-the-art AI diplomats (Bakhtin et al., 2022).
@@
Our contributions can be summarized in three-fold: 1) We introduced a new paradigm for building AI diplomacy agents, compared to previous work (Fig. 1). The agent can self-evolve by generating experience via self-play games, without any task-specific human data. 2) We demonstrate the superior performance of our agent playing against the SOTA method, e.g., Cicero Bakhtin et al. (2022), that relies on a large-scale human demonstration for training. 3) We further analyze the effectiveness of each module in our agent and the generalization of our agent in adopting different LLMs, such as GPT4.0 and Llamma 3.
@@ 2 Related work
AI Diplomacy. The game involves seven players controlling different powers in Europe. In each turn, players can negotiate for cooperation before making moves to take as much supply centers as they can. Apparently, this challenging strategy task requires both complex negotiation skills and superior planning capability for player agents to achieve final victory. So far, most AI research on this task remain focused on the planning strategies (a.k.a. No-Press Diplomacy where no communication channels are allowed). The setting remains challenging considering its enormous action space of $110^{2}$ 1 to $410^{6}$ 4 per turn (compared with Chess, which has much fewer than 100 actions per turn). No wonder existing efforts rely on human data to play the game. Among the methods, one typical research is DipNet (Paquette et al., 2019) which uses supervised and reinforcement learning. Based on DipNet, BRPI (Anthony et al., 2020), SearchBot (Gray et al., 2020), DORA (Bakhtin et al., 2021), and KL-Regularized search (Diplodocus) (Jacob et al., 2022) were conducted. Until very recently, research has also emerged for the full-setting of Diplomacy, or Press Diplomacy where players are allowed to communicate with each before making their moves in each turn. Such studies (De Jonge and Sierra, 2017)(Bakhtin et al., 2022)(Jaidka et al., 2024)(Kramár et al., 2022) mainly benefit from the recent thriving language models. Specifically, notable advancements include policy iteration methods from DeepMind and Facebook AI Research’s equilibrium search agent (Jaidka et al., 2024). However, Deepmind propose to learn negotiation agents based on predefined contracts/protocols (Kramár et al., 2022). And Meta AI’s work, instead of one unified architecture, Cicero (Bakhtin et al., 2022) integrates a language model for negotiation and an RL model for planning respectively. Such separately trained models make it inconvenient for agents’ continual evolution. What’s more, like no-press methods, these approaches heavily rely on human player data for agent training. Unlike these approaches, this paper delves into solving the negotiation and planning in one single self-evolving LLM-based agent model, without any pre-collected human expert training data.
@@
LLM-based Agents. With the emergence and growth of large language models (LLM), there is a growing trend in utilizing LLMs as fundamental controllers for autonomous agents(Wang et al., 2024c). One wide application genre is LLM-based answering engines, which merely cover the negotiation aspects of Diplomacy. Such systems include HuggingGPT (Shen et al., 2023), GPT4Tools (Yang et al., 2023b) and ToT Yao et al. (2023), etc. They leverage LLMs to manage Al models, use tools, implement policy iteration, and enhance problem-solving across various tasks. Related work including AutoGPT, AgentGPT, BabyAGl (Talebirad and Nadiri, 2023), Toolformer (Schick et al., 2023), and Visual ChatGPT aim to improve LLM capabilities in task automation and tool usage. And Reflexion, a framework that improves LLMs through linguistic feedback and episodic memory (Zhang et al., 2023), facilitating better decision-making across diverse tasks is proposed. Besides (Wang et al., 2023b)(Wang et al., 2023a)(Wang et al., 2023c)(Zhu et al., 2023)(Yan et al., 2023) apply LLM agents to the complex planning tasks in the well-known open-world game Minecraft(Fan et al., 2022). Unlike these LLM-based agents which only focus on the negotiation/planning aspect, the proposed approach involves multiple self-evolving schemes to handle both of them simultaneously.
@@ 3 Problem Statement
The Diplomacy game Wikipedia (2024); Calhamer (1974) is set in pre-World War I Europe and involves each player (agent) representing one of the seven Great Powers of Europe, such as Germany, France, England, Italy, Austria-Hungary, Russia, and Turkey. Each player has a set of military units, including armies and fleets, which they can move and use to capture other supply centers. The ultimate goal for the agent is to control a majority of the total supply centers on the board by the end of the game’s Fall phase. It’s important to note that it is not won by eliminating other players or their units; it is won by controlling the requisite number of supply centers. This often involves forming and breaking alliances, negotiating, and sometimes betraying other players to achieve one’s own goals.
In each turn,Wei et al. (2022); Wang et al. (2022) the agent $i$ gets the current state $s_{t}\in S$, the actions of other players from the previous turn $\vec{a}^{-i}_{t-1}$, and the messages $\vec{m}^{-i,i}_{t}$ from other players during this turn’s negotiations. The state $s_{t}$ for the environment includes the ownership of each territory on the map by a particular country and where the armies of each country are located. Based on this information, the agent needs to engage in negotiations with other players, sending messages $\vec{m}^{i,-i}_{t}$ to chat with other players, and then take the actions $a^{i}_{t}$ in this turn. The possible actions an agent can take $a^{i}_{t}\in A$ are commands to the armies, such as moving into an adjacent territory, supporting another unit, or holding a position. Actions can also include diplomatic moves, such as proposing or withdrawing from an alliance, although these are less formalized in the game mechanics.Paquette et al. (2019); Hill (2014)
@@ 4 Self-Evolving LLM-based Diplomat
We have constructed a comprehensive framework with modules for memory management, social reasoning, strategic planning, negotiation, decision-making, memory update, and self-evolving to fully leverage the capabilities of LLMs. Richelieu starts by setting up with map details, game rules, domain knowledge, and the long-term goal.Zhang et al. (2022); Wei et al. (2022); Wang et al. (2022) At each turn, the agent will run in the following steps: 1) Social Reasoning: First of all, the agent undergoes a comprehensive analysis of the game state $s_{t}$ to build the social belief, including the intention of other players and their relationship $\vec{\phi_{t}}\in\Phi^{n}$.Zhang et al. (2024c); Gürcan (2024) 2) Planner with Reflection: Then, the agent proposes sub-goals $\chi^{i}_{t}\in X$ that is strategically aligned with the long-term goals $\Upsilon$, with the social belief and refining the proposed goal with experience $\vec{\eta_{t}}\in H^{m}$ abstract from the memory $M$ via self-reflection.Wang et al. (2024b, d) 3) Negotiator: To achieve the sub-goals, the negotiator will start a dialogue session with some players, and evaluate their trueness $\vec{\psi}^{-i}_{t}$ by referring to their words $\vec{m}^{-i,i}_{t}$, the current state $s_{t}$, their sincerity $\vec{\gamma}^{-i}_{t}$ and the experience $\vec{\xi_{t}}$Abdelnabi et al. (2023); Bianchi et al. (2024) 4) Actor: After negotiation, the actor decides its course of action $a^{i}_{t}$, based on the sub-goal $\chi^{i}_{t}$ and updated social state $s_{t+1}$, marking the end of that turn. 5) Memory Management: The state of the current turn $s_{t}$, the content of negotiations $\vec{m_{t}}$, the actions taken by all players $\vec{a_{t}}\in A^{n}$, and the sub-goals set forth $\chi^{i}_{t}$ are all logged within the memory as $\mu\in M$. This logged data serves as a historical experience, guiding Richelieu’s subsequent actions in future turns Hatalis et al. (2023); Zhang et al. (2024e). 6) Self-evolution: The agent’s evolution is highly dependent on the diversity of experiences stored in its memory. As this diversity grows, so does the agent’s capability. Without human demonstrations, we employ multi-agent self-play games, i.e., our agents respectively control all the countries to simulate and acquire diverse experiences for self-evolving. Notably, the agent can further evolve during testing to adapt to different players.
@@ 4.1 Social Reasoning
There are no permanent enemies, no permanent allies. The relationship among countries is dynamically changing upon the evolving global state. However, it is difficult to determine the appropriate allies and enemies with partial observation. For example, there is uncertainty about the intentions of potential allies, which could lead to betrayal at pivotal moments. Consequently, we need to identify the intention and relationship of the current state by social reasoning to shape the social belief Zhang et al. (2024c); Gürcan (2024).
1) Modeling Relationship: Before setting sub-goals, Richelieu evaluates its relations with other players, identifying enemies such as aggressive nations, vulnerable neighbors for expansion, and those with long-term potential threats. It also seeks out potential allies to counter these threats.Sun et al. (2024); Zhang et al. (2024d) Simultaneously, Richelieu also tries to identify potential allies that could be instrumental in countering these adversaries. By isolating the analysis of inter-player relationships as a discrete element, Richelieu strategically exploits the actions of other players in subsequent stages of the game to reach its goals. 2) Inferring Intention: The social belief is also used by the planner, ensuring that Richelieu’s sub-goals are formulated with a comprehensive consideration of the behaviors and intentions of other intelligent agents within the game. Richelieu’s sub-goals will particularly emphasize on those who are identified as potential adversaries or allies, fostering more effective collaboration with potential allies and participation in strategic opposition against adversaries. Furthermore, the insights gleaned from this analysis are instrumental in the subsequent negotiation phases. They are employed to assess the authenticity of the statements made by other players, as well as to aid Richelieu in reaching cooperative agreements.de Zarzà et al. (2023); He et al. (2024)
@@ 4.2 Strategic Planner with Reflection
The strategic planner specifies the sub-goals, which serves as an intermediary between immediate actions and the overarching goal of securing victory in the game. That is because we observe that LLMs are often characterized by their propensity to prioritize short-term gains in decision-making processes, with a notable deficiency in incorporating the future into their strategic calculations. Renze and Guven (2024); Zhang et al. (2024b)For example, it is common for a non-neighboring country to become too powerful. Formally, $\vec{\chi_{t}}\leftarrow SR(s_{t},\vec{\phi_{t}},\Upsilon)$ where $\vec{\chi_{t}}=(\chi^{i}_{t},\chi^{1}_{t},\ldots,\chi^{n}_{t})$ represents the proposed sub-goals and other players’ intention that we inferred, $\vec{\phi_{t}}\in\Phi^{n}$ represents the inferred relationship on the social belief. These goals may encompass a range of tactical considerations, such as the containment of a formidable rival’s advancement or the strategic expansion in a particular direction to consolidate power.
Reflection with Memory. We further develop a reflection mechanism to enhance the rationality and effectiveness of our agent’s sub-goals in achieving long-term goals.Liu et al. (2024) This reflection mechanism relies on the past experiences to critique and enhance proposed sub-goals. We employ a similarity-based function to find relevant historical experiences that match the current game state from its memory. This function considers two factors: goal similarity and state similarity, to select the most comparable experiences. The process can be written as: $\vec{\eta_{t}}\leftarrow h(s_{t},\chi^{i}_{t},M)$, where $\vec{\eta_{t}}\in H^{m}$. In practice, considering the limited context windows of LLM, we retrieve the most analogous experiences from the memory based on these metrics. Experiences with high evaluative scores reinforce successful strategies and support the continuity of existing sub-goals. On the other hand, lower scores indicate areas that need improvement and prompt the necessary adjustments. As our agent, Richelieu, undergoes more training sessions, its reflection abilities improve. The growing pool of historical experiences consistently enhances its performance.